. A ball is launched horizontally from the edge of a table. It strikes the floor. The horizontal distance between the point on the floor directly beneath the edge of the table and the point on the floor where the ball strikes is 3.23 m. The vertical distance between the floor and the edge of the table is 1.14 m. Find (a) the time interval (s) from when the ball was launched to when the ball struck the floor and (b) the initial speed (m/s) of the ball. Next, the ball is launched from the floor at an angle 56.1 ◦ above the horizontal with the same initial speed. It again strikes the floor. Find (c) the time interval (s) from when the ball was launched to when the ball strikes the floor and (d) the horizontal distance (m) from the point where the ball was launched to the point where the ball strikes the floor.
Answer #1

a) y = vyt - 1/2at2

-1.14 = - 4.9t2

t = 0.48234 sec

------------------------------------------------------------------------------

b) x = vxt

vx = x/t

vx = 3.23 / 0.48234

vx = 6.696 m/s

--------------------------------------------------------------------------------------

now when launched at an angle,

time = 2vsin(theta) / g

time = 2*6.696*sin56.1 / 9.8

time = 1.134 sec

-----------------------------------------------------------------

range = horizontal distance

R = 6.6962sin 112.2 / 9.8

R = 4.235 m